solve by substitution/elemination
{x-2y+3z=2} {-x-5y+2z=-11} {2x-y-4z=0}

Answer by mananth(16946)

x-2y+3z= 2-------------- 1
-x-5y+2z =-11--------------2
2x-y-4z= 0-------------- 3
consider equation 1 &2 Eliminate y
Multiply 1 by -5
Multiply 2 by 2
we get
-5x+10y-15z=-10
-2x-10y+ 4z=-22
Add the two
-7x+-11z =-32-------------4
consider equation 2 & 3 Eliminate y
Multiply 2 by -1
Multiply 3 by 5
we get
x+5y-2z= 11
10x-5y-20z=0
Add the two
11x-22z= 11-------------5
Consider (4) & (5)Eliminate x
Multiply 4 by -11
Multiply (5) by 7
we get
-77x+121z=352
77x-154z =77
Add the two
154x-33z =429
/-33
z=-13
Plug the value of z in 5
11x+286 =11
11x=-275
x=-25
plug value of x & z in 1
-25-2y-39=2
-2y=25+39+2
y=66